Find the length of a rectangular lot with a perimeter of 82 m if the length is 5 m more than the width.

Answer by Mr.Math(29)

Let the length=a, and width=b, then, by the condition of the problem:
a = b+5
P = 2(a+b)
a = b+5
82 = 2*(a+b)
a = b+5
82 = 2* (b+5+b)
a = b+5
82 = 4b + 10
b = 18 => a=18+5=23
So, length= 23m